# ADVANCING MATHEMATICS RESEARCH WITH LARGE LANGUAGE MODELS

### v1

@@ 15.1. Design your outputs
In LLMs, there is a ‘system instruction’ feature which can be used to influence the type of output. You can enter a ‘pre-prompt’ with your own personal profile that governs the model’s behavior for the entire chat session.
@@ 15.2. Changing the ‘knobs’
System settings in an LLM can also be manually changed in order to make mathematical proofs more rigorous and less random.
For mathematical rigor, the most important setting is temperature’. This controls the randomness of the output by reshaping the probability distribution. A high temperature (such as 0.9) encourages more diverse and novel responses. A low temperature makes the output more deterministic. The recommended setting for math in Gemini is to lower the temperature to 0.2 or 0.1.
@@ 15.3. Advanced: system prompts
‘Back-end’ engineering, such as writing system prompts in an LLM gives more direct control and can impact the amount of detail and rigor in a proof.
16. Hybrid systems: LLMs with Computer Algebra Systems or formal proof assistants
Because LLMs are not substitutes for formal proof or verified computation, their outputs require rigorous human scrutiny. The primary way to offset this limitation is to use them in conjunction with formal tools like Computer Algebra Systems or formal proof assistants such as Lean. For example, ChatGPT can integrate with WolframAlpha. The user can explicitly prompt ChatGPT in the Wolfram chat window, or allow it to decide when to outsource a computation.
@@ 16. Hybrid systems: LLMs with Computer Algebra Systems or formal proof assistants
However, writing Lean code is notoriously difficult, requiring not just programming skill but the precise formalization of abstract mathematical ideas in Type Theory. In this framework, even small logical gaps or unstated assumptions must be made explicit and rigorously verified. Precise use of syntax is required. One of the benefits of the Lean environment is Lean’s Mathlib: a vast, open-source digital encyclopedia of formally verified mathematics.
This is where LLMs can potentially play a transformative role. Research labs are actively working to bridge this difficulty by using LLMs to write Lean code and accept inputs in natural language, making formal verification more accessible.
17. Math research alongside LLMs and Lean
This is a brief description of joint research with Ayush Khaitan and Scott H. Murray alongside LLMs and Lean [CKM]. Results have been verified using Lean 4.
The project was to study an infinite dimensional analog $S(\mathfrak{m})$ of the Steinberg group, given by generators and relations, associated to the Monster Lie algebra $\mathfrak{m}$. This is an infinite dimensional Lie algebra which is an example of a Borcherds algebra, also called a generalized Kac–Moody algebra.
We proposed a presentation for the group and used LLMs is to assist in determining properties of $S(\mathfrak{m})$.
The work was done using a combination of usual research techniques, laborious use of context prompting, chain-of-thought prompting, and iterative prompting of LLMs, particularly Gemini and ChatGPT, verification and determining errors in the output and reprompting.
17.1. New results suggested by Gemini
During a long chat with Gemini, the model suggested the following:
Working with the group presentation indicates that the group $S(\mathfrak{m})$ is a perfect group. That is, $S(\mathfrak{m})=(S(\mathfrak{m}),S(\mathfrak{m}))$ is equal to its commutator subgroup. This property holds for certain Steinberg groups of finite dimensional simple Lie algebras.
Since the setting of finite dimensional Steinberg groups is vastly different, this had not seemed like a plausible research direction before Gemini’s suggestion.
The usual method for proving that Steinberg groups of finite dimensional simple Lie algebras are perfect involves showing that each generator is contained in the commutator subgroup. The result turned out to be true for a larger group $G(\mathfrak{m})$ constructed in [CJM], containing toral elements.
Gemini also aided in completing the highly technical proof and suggested that the group $G(\mathfrak{m})$ behaves like a simple group. This is a plausible suggestion which was already being considered.
17.2. Challenging calculations
Let $L$ be a $\mathbb{Z}\oplus\mathbb{Z}$-graded free Lie algebra on a specified set of generators.
Two different LLMs were used to obtain recursive formulas for the dimensions of the homogeneous components $L_{(m,n)}$ of $L$, for $m,n>0$.
Given $(m,n)$ with $m>0,n>0$, let $P(m,n)$ be the set of all partitions of the ordered pair $(m,n)$ into a sum of the form ${\sum}_{r,s}a_{rs}(r,s)$ such that $a_{rs},r,s\in\mathbb{N}_{\geq 1}$, ${\sum}_{r,s}a_{rs}\cdot r=m$, and ${\sum}_{r,s}a_{rs}\cdot s=n$.
For example $P(3,2)=\{(3,2),(1,1)+(2,1)\}$ where $a_{3,2}(3,2)=1(3,2)$ and $a_{1,1}(1,1)+a_{2,1}(2,1)=1(1,1)+1(2,1)$.
Given the string of coefficients $(a_{r,s})$, and a partition $a\in P(m,n)$, [JurJPAA1998] and [JLW1995] gave the formula:
$d(i,j)={\sum}_{k\in{\mathbb{Z}}_{+}\atop k(m,n)=(i,j)}{1\over k}\mu(k){\sum}_{a\in P(m,n)}{(\sum a_{rs}-1)!\over\prod a_{rs}!}\prod n(r,s)^{a_{rs}}$
where $n(i,j)$ is the number of generators of the algebra $L$ of degree $(i,j)$ and $\mu(k)$ is the Möbius function.
Using a staged prompting method, Gemini and ChatGPT/WolframAlpha produced the same symbolic expressions for $d(2,j)$ in terms of $n(r,s)$, for $j$ from 1 to 50, by first computing the partitions and then the $d(i,j)$. Both LLMs also gave a recursive strategy for the calculation of $d(i,j)$ in general.

### v2

@@ 16.1. Design your outputs
In generative AI models, there is a ‘system instruction’ feature which can be used to influence the type of output. You can enter a ‘pre-prompt’ with your own personal profile that governs the model’s behavior for the entire chat session.
@@ 16.2. Changing the ‘knobs’
System settings in certain generative AI models can also be manually changed in order to make mathematical proofs more rigorous and less random.
For mathematical rigor, the most important setting is ‘temperature’. This controls the randomness of the output by reshaping the probability distribution. For research questions and formulating conjectures, a high temperature (such as 0.9) is preferable, as it encourages more diverse and novel responses. A low temperature (such as 0.2 or 0.1) makes the output more deterministic.
Table 2. Optimized settings in Google AI studio for research in mathematics
Temperature Medium-High ($0.7$ - $1.0$) A higher temperature encourages exploration.
Top P High ($0.95$ - $1.0$) Allows the model to consider a wider, more diverse set of ‘next words’.
Thinking mode Advanced Research problems require the deepest level of reasoning.
Set thinking budget Maximum Gives the model time to explore the ‘search space’.
Code execution On Crucial for running checks with Python tools such as SymPy.
Grounding (Google Search) On Allows the model to search outside its training data.
URL Context On Allows you to give the model specific URLs as context prompts.
Structured output Off Allows unstructured brainstorming.
Function calling On Connect to external research tools such as a university library API or a computational algebra system like SageMath or WolframAlpha.
Media resolution High Essential for correct reading of symbols in handwritten notes or complicated diagrams.
Output Length Maximum Allows for an unencumbered output.
@@ 16.3. Advanced: system prompts
‘Back-end’ engineering, such as writing system prompts in a generative AI model gives more direct control and can impact the amount of detail and rigor in a proof.
17. Hybrid systems: generative AI models with Computer Algebra Systems or formal proof assistants
Because generative AI models are not substitutes for formal proof or verified computation, their outputs require rigorous human scrutiny. The primary way to offset this limitation is to use them in conjunction with formal tools like Computer Algebra Systems or formal proof assistants such as Lean. For example, the professional version of ChatGPT can integrate with WolframAlpha. The user can explicitly prompt ChatGPT in the Wolfram chat window, or allow it to decide when to outsource a computation.
It is an active area of research to develop tools that link generative AI models with Computer Algebra Systems. One such tool is Google DeepMind’s AlphaGeometry. See also [KG2025].
@@ 17. Hybrid systems: generative AI models with Computer Algebra Systems or formal proof assistants
However, writing Lean code is notoriously difficult, requiring not just programming skill but the precise formalization of abstract mathematical ideas in Type Theory. In this framework, even small logical gaps or unstated assumptions must be made explicit and rigorously verified. Precise use of syntax is required. One of the benefits of the Lean environment is Mathlib: a vast, open-source digital encyclopedia of formally verified mathematics.
Research labs are actively working on ways to make the task of generating Lean code easier: from using generative AI models to write Lean code, to developing interfaces that accept inputs in natural language.
18. Should I collaborate with a generative AI model?
As we have discussed, generative AI, on its own, should not be viewed as an oracle or an authoritative source for mathematical output. Verification of output is a necessary part of using an AI model for mathematics. But it can be used as a creative collaborator that provides ideas, perspective and potential new directions. When combined with a Computer Algebra System, it can reliably perform laborious tasks and calculations. When paired with formal verification tools, it transforms the research process itself.
The author of this work now has several LLM+CAS+Lean-assisted research projects in infinite dimensional algebra and group theory. The use of generative AI models has been particularly revealing. There have been several instances where AI models have suggested novel research directions that had initially been dismissed as implausible or irrelevant. Yet, these suggestions turned out to yield fruitful and unexpected paths forward.